OpenStudy (anonymous):

What is the solution of the graphs of two lines: -6x-10y=-12 3x+5y=-9

OpenStudy (anonymous):

plug into a matrix and solve

OpenStudy (anonymous):

What you can do is make theme equal to eachother

OpenStudy (tkhunny):

-6x-10y=-12 3x+5y=-9 Multiply the second by 2 -6x-10y=-12 6x+10y=-18 I see the solution, now. Do you?

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

No

OpenStudy (tkhunny):

Add the two equations. What happens?

OpenStudy (anonymous):

Draw the two lines on a graph and find where they intersect

OpenStudy (anonymous):

Yeah but how do I draw the two lines

OpenStudy (tkhunny):

Seriously, add the two equations. If you MUST graph them, the Slope-Intercept form is often useful. Put them in that form.